# MLLM-Tool: A Multimodal Large Language Model For Tool Agent Learning

### v2

@@ 6.2 Data processing
Vicuna-7B (3epochs) 77.35 4.13 100.00
Llama-7B (3epochs) 76.64 2.80 100.00
Llama2-7B (3epochs) 76.39 5.13 100.00
Llama2-Chat-7B (3epochs) 76.18 4.92 100.00
Vicuna-7B (5epochs) 83.15 1.63 100.00
Llama-7B (5epochs) 84.48 0.58 100.00
Llama2-7B (5epochs) 80.68 1.96 100.00
Llama2-Chat-7B (5epochs) 83.60 1.34 100.00
Vicuna-13B (3epochs) 74.34 3.63 100.00
Llama-13B (3epochs) 75.18 3.21 100.00
Llama2-13B (3epochs) 64.71 3.38 100.00
Llama2-Chat-13B (3epochs) 77.51 3.38 100.00
Vicuna-13B (5epochs) 80.85 1.63 100.00
Llama-13B (5epochs) 81.44 2.00 100.00
Llama2-13B (5epochs) 67.58 1.34 100.00
Llama2-Chat-13B (5epochs) 81.35 1.42 100.00
Vicuna-13B (10epochs) 88.19 0.17 100.00
Llama-13B (10epochs) 87.86 0.17 100.00
Llama2-13B (10epochs) 81.81 0.17 100.00
Llama2-Chat-13B (10epochs) 87.78 0.08 100.00
Table 5: Experimental results of MLLM-Tool. We present the performance of eight large language models with different configurations, and we report the performance at different epochs of training to reflect the changing trend.
[…]
Vicuna-7B (3epochs) 60.92 0.00 64.71 3.92 97.06 0.00 87.86 0.00 97.92 0.00 82.74 0.38 74.89 3.49
Llama-7B (3epochs) 54.02 1.15 66.67 0.00 47.06 0.00 87.54 0.00 100.00 0.00 78.61 0.19 76.98 5.26
Llama2-7B (3epochs) 58.62 0.00 64.71 0.00 61.76 0.00 87.86 0.32 97.92 0.00 80.11 0.19 73.50 4.02
Llama2-Chat-7B (3epochs) 66.67 0.00 74.51 0.00 79.41 0.00 88.50 0.00 100.00 0.00 84.05 0.00 73.93 6.33
Vicuna-7B (5epochs) 83.91 2.30 72.55 0.00 100.00 0.00 94.25 0.00 100.00 0.00 91.37 0.94 80.79 1.98
Llama-7B (5epochs) 75.86 0.00 66.67 0.00 73.53 0.00 95.21 0.00 100.00 0.00 88.37 0.00 83.37 0.75
Llama2-7B (5epochs) 66.67 0.00 66.67 0.00 94.12 0.00 92.97 0.00 100.00 0.00 86.87 0.00 78.92 2.52
Llama2-Chat-7B (5epochs) 87.36 0.00 74.51 0.00 97.06 0.00 95.85 0.00 100.00 0.00 92.87 0.00 80.95 1.72
Vicuna-13B (3epochs) 56.32 0.00 64.71 0.00 82.35 0.00 88.82 0.00 100.00 0.00 81.80 0.00 73.28 4.13
Llama-13B (3epochs) 24.14 0.00 66.67 0.00 44.12 11.76 89.46 0.00 100.00 0.00 74.67 0.75 74.25 4.45
Llama2-13B (3epochs) 59.77 0.00 76.47 0.00 58.82 0.00 90.10 0.00 97.92 0.00 82.55 0.00 76.07 4.35
Llama2-Chat-13B (3epochs) 39.08 0.00 68.63 0.00 38.24 0.00 26.20 1.60 97.92 0.00 39.59 0.94 71.89 4.08
Vicuna-13B (5epochs) 82.76 0.00 68.63 7.84 97.06 0.00 93.61 0.32 100.00 0.00 90.24 0.94 78.92 2.31
Llama-13B (5epochs) 48.28 0.00 62.75 0.00 100.00 0.00 88.50 0.96 100.00 0.00 81.24 0.56 80.74 1.93
Llama2-13B (5epochs) 74.71 0.00 60.78 0.00 91.18 0.00 88.82 0.00 100.00 0.00 84.99 0.00 80.31 1.82
Llama2-Chat-13B (5epochs) 33.33 0.00 58.82 0.00 41.18 0.00 31.63 0.00 87.50 0.00 40.15 0.00 75.43 1.72
Vicuna-13B (10epochs) 88.51 0.00 66.67 0.00 97.06 0.00 97.76 0.00 100.00 0.00 93.43 0.00 86.70 0.21
Llama-13B (10epochs) 89.66 0.00 76.47 0.00 100.00 0.00 97.12 0.00 100.00 0.00 94.37 0.00 86.00 0.21
Llama2-13B (10epochs) 64.37 0.00 60.78 0.00 73.53 0.00 73.48 0.32 100.00 0.00 73.17 0.19 84.28 0.16
Llama2-Chat-13B (10epochs) 90.80 0.00 78.43 0.00 97.06 0.00 97.44 0.00 100.00 0.00 94.75 0.00 85.78 0.11
Table 6: The performance on testing subsets divided by ambiguity type.
6.3 Experimental Results
We finetune four open-source LLMs, Llama, Llama2, Llama2-Chat, and Vicuna, each with 7B and 13B two variants, on the instruction-solution data pairs. As illustrated in Table 5, we report the format accuracy, the accuracy, and the hallucination rate, which are defined in Table 5 in detail. Due to the limitation of computing resources, the experiment results for recall metric, we focused on the two models, Llama2-7B and Llama2-Chat-7B, and set the inference times $K$ 10. The results are shown in Table 7.
[…]
Llama2-7B (3 epochs) 55.85
Llama2-7B (5 epochs) 62.13
Llama2-Chat-7B (3 epochs) 55.23
Llama2-Chat-7B (5 epochs) 61.47
Table 7: The recall performance of two 7B models.
We can derive some findings from the results:
1. For the 7B models, Llama-7B performs the best performance both on accuracy and hallucination rate, with 84.48% and 0.58% respectively when trained with 5 epochs. As we increase the training epochs, we can observe that there is a consistent improvement in both accuracy (+ 6.34%) and the reduction of hallucination rates (- 2.87%) averagely compared to that with 3 epochs. As for 13B models, when training for 5 epochs, the performance is not as good as the corresponding 7B model. But if increasing to 10 epochs, we can observe a substantial improvement, and Vicuna-13B achieves 88.19% accuracy. Meanwhile, all the 13B models’ hallucination rate drop to within 0.17% at 10 epochs.
2. The format maintains a hundred percent accuracy across all models, demonstrating LLMs have a high understanding of structural rules, which is crucial for tasks that demand strict formatting.
3. The hallucination rate generally remains low, especially when increasing the number of epochs, which we speculate is due to the relative simplicity of our dataset. Since most of our data provides a clear context and our extracted description does not involve complex language structures so that the model is more likely to produce responses within the whole corpus.
4. While recall gradually increases with the number of epochs, the overall recall values remain modest. We hypothesize this is due to an abundance of selectable APIs for some queries. Despite limiting identical function APIs, in cases like image classification, common object labels appearing across multiple datasets can accumulate suitable APIs, potentially exceeding ten options. Therefore, even with ten inference times, it’s impossible to cover all possibilities.
Overall, the results affirm that our system MLLM-Tool effectively brings out the tool-use potential in LLMs and enhances the LLMs to deal with multi-modal instruction inputs.
Model Video Audio Image Text
[…]
Vicuna-7B (3epochs) 76.19 4.76 88.41 3.05 79.87 1.18 73.62 3.58
Llama-7B (3epochs) 85.71 0.00 92.68 0.00 75.66 1.18 76.38 6.20
Llama2-7B (3epochs) 80.95 0.00 89.63 0.61 76.97 1.05 72.18 4.61
Llama2-Chat-7B (3epochs) 76.19 0.00 87.20 0.00 79.87 0.66 73.00 7.78
Vicuna-7B (5epochs) 71.43 0.00 93.29 0.00 88.95 0.53 79.13 2.41
Llama-7B (5epochs) 76.19 0.00 89.02 0.00 87.24 0.26 82.64 0.83
Llama2-7B (5epochs) 80.95 0.00 93.29 0.00 83.82 0.39 77.62 3.03
Llama2-Chat-7B (5epochs) 80.95 0.00 92.07 0.00 86.45 1.84 81.20 1.24
Vicuna-13B (3epochs) 85.71 0.00 89.63 1.83 76.71 1.84 72.59 4.13
Llama-13B (3epochs) 80.95 0.00 92.07 0.00 72.50 1.45 73.21 5.23
Llama2-13B (3epochs) 76.19 0.00 89.63 0.00 78.16 1.32 75.83 4.89
Llama2-Chat-13B (3epochs) 85.71 0.00 90.24 0.00 44.61 1.32 72.04 4.89
Vicuna-13B (5epochs) 80.95 0.00 89.63 0.00 86.18 1.97 78.03 2.27
Llama-13B (5epochs) 76.19 0.00 90.85 0.00 79.87 0.92 80.30 2.20
Llama2-13B (5epochs) 71.43 0.00 92.07 0.00 82.11 0.13 79.89 2.27
Llama2-Chat-13B (5epochs) 76.19 0.00 85.37 0.00 44.21 0.53 77.69 1.93
Vicuna-13B (10epochs) 76.19 0.00 92.07 0.00 90.66 0.13 86.64 0.21
Llama-13B (10epochs) 71.43 0.00 92.07 0.00 92.76 0.00 85.06 0.28
Llama2-13B (10epochs) 85.71 0.00 89.02 0.00 71.84 0.13 86.16 0.21
Llama2-Chat-13B (10epochs) 80.95 0.00 89.63 0.00 90.53 0.00 86.23 0.14
Table 8: The performance on testing subsets divided by input modality.
6.4 Results Analysis
To verify the effectiveness of our system when encountering different complex cases, we conduct experiments on all types of testing subsets defined by three previously established criteria in Sec. 5, assessing the impact of different characteristics of data on our model’s performance and accuracy. All the 7B models are trained in 5 epochs and 13B models are trained in 10 epochs by default for a fair comparison.
Ambiguity types. In Table 6, we provide the results for five types of ambiguity, and the comparison between non-ambiguity and ambiguity cases. We find that the ambiguity types I (Domains) and II (Categories) perform worst in ambiguity handling. In contrast, other categories excel, especially type V (Language) achieves 100% accuracy when trained 5 epochs for 7B or 10 epochs for 13B models. Intuitively, the exceptional performance of type III (Quality) and type IV (Conditions) may stem from significant visual differences and clear classification boundaries. In contrast, for type I and type II, particularly the latter, the challenges arise due to varying granularity in datasets. The model must not only understand the multimodal input content but also discern the most appropriate classification granularity for the task at hand, complicating the choice of API. For the overall performance, we find that performance without ambiguity surprisingly falls short compared to that with ambiguity. It indirectly highlights that text descriptions cannot substitute for multimodal information, even when the language is exceptionally detailed. This underscores the necessity for LLMs to process the capability to perceive visual- and auditory- instructions, demonstrating the significance of our work.
Model one-to-one one-to-many
[…]
Vicuna-7B (3epochs) 51.75 6.71 86.62 1.23
Llama-7B (3epochs) 56.12 6.85 85.86 3.04
Llama2-7B (3epochs) 47.67 6.56 85.91 1.81
Llama2-Chat-7B (3epochs) 57.29 6.56 83.75 4.27
Vicuna-7B (5epochs) 66.03 3.35 90.01 0.94
Llama-7B (5epochs) 69.24 1.31 90.59 0.29
Llama2-7B (5epochs) 63.12 3.64 87.73 1.29
Llama2-Chat-7B (5epochs) 70.99 2.33 88.66 0.94
Vicuna-13B (3epochs) 47.96 7.14 86.09 1.64
Llama-13B (3epochs) 47.23 5.69 85.21 2.81
Llama2-13B (3epochs) 51.90 8.16 87.78 1.46
Llama2-Chat-13B (3epochs) 41.84 7.00 73.87 1.93
Vicuna-13B (5epochs) 61.95 4.52 89.25 0.99
Llama-13B (5epochs) 61.81 3.35 88.49 0.94
Llama2-13B (5epochs) 62.97 3.06 88.72 0.76
Llama2-Chat-13B (5epochs) 50.58 3.35 74.40 0.53
Vicuna-13B (10epochs) 77.26 0.58 92.58 0.00
Llama-13B (10epochs) 76.68 0.58 92.34 0.00
Llama2-13B (10epochs) 69.24 0.44 86.85 0.06
Llama2-Chat-13B (10epochs) 76.38 0.29 92.34 0.00
Table 9: The performance on testing subsets divided by API option number.
Multiple options. We divide the testing set into two to distinguish whether prompts can be linked to multiple API options (one-to-one or one-to-many), and the corresponding performances are shown in Table 9. Surprisingly, we discover that one-to-one performance does not surpass, and is significantly behind, the scenario with one-to-many cases, lagging by an average of 21.90% across four 7B models at 5 epochs and 16.14% across four 13B models at 10 epochs . Besides, in terms of the rate of hallucinations, the one-to-one demonstrates a relatively higher incidence. Our analysis suggests that one-to-one pairs may lead to overfitting, where the model becomes overly tuned to specific pairs, hindering its ability to generalize to slightly different queries or contexts. The higher rate of hallucinations also verifies the above speculation.
Individual modality. We further investigate the model’s perception and understanding of instructions with the specific sole modality and report the results in Table 8. Notably, the audio modality stands out, achieving an impressive performance among all models, reaching as high as 93.29% on both the Llama2-Chat-7B and Vicuna-7B model. On the other hand, we notice that pure text-based performance is not particularly satisfactory, falling short in both accuracy and hallucination rate. As for the video modal, it has a greater performance variance than other modals. Interestingly, We find that as the number of epochs increased, the performance of Vicuna-7B drops significantly. A similar situation also occurs on the several other models. We speculate that this may be related to the complexity and scarcity of video data. Since our video task mostly involves video understanding, but with the epochs increasing, the model might start to emphasize the detailed information, which can impact the performance when extending to new unseen data.
Figure 5: Qualitative example of MLLM-Tool on text ambiguity case. MLLM-Tool firstly receives the text instruction and perceives information from the input image to determine the corresponding task, select the appropriate API that meets the requirements from the API corpus, and then call the API to execute, and finally get the quality-improved image. Zoom in for the best view.
6.5 Visualization
We visualize qualitative result of MLLM-Tool, to verify its effectiveness. As shown in Figure 5, our system is asked to improve the quality of the input image without expaining the reason for the poor image quality, and MLLM-Tool is capable of perceiving the multi-modal information and choosing the correct API by combining the instruction and visual information, which proves that our MLLM-Tool can effectively accomplish the users’ needs.
7 Conclusion
In this work, we introduce how to empower the tool-augmented LLMs with perceiving visual- or audio-grounded instructions’ information. We collect a benchmark, ToolMMBench, from the HuggingFace Platform, which covers 932 high-quality machine learning related APIs. The finetuning of this dataset on various popular LLMs has yielded promising results, demonstrating that MLLM-Tool can accurately recommend appropriate tools for multi-modal instructions.
[…]
Supplementary Material
In this supplementary material, we list i) dataset details, including selection criteria, prompt templates, data distribution; ii) our proposed categorization system.
8 Dataset Details
In this section, we elaborate on the details of dataset construction and the distribution of training and testing data. Specifically, Figure 8.1 visualizes the entire dataset construction process, while Sec. 8.2 - Sec. 8.4 detail each step, including API selection criteria, API functional boundaries criteria, and prompt construction. Sec. 8.5 visualizes some data distributions of the training set and testing set. We emphasize that the collection and processing of this dataset involved over 2000 person-hours, ensuring its reliability for advanced computer vision research and applications by meticulous attention to detail.
Figure 6: The process of dataset construction. It mainly consists of three stages. The first step is collection and preprocessing, including crawling the original API from HuggingFace, filtering and re-categorizing the API according to tasks. The second step is the generation of instruction-answer pairs, including using GPT-4 to generate instructions based on API functions, determining whether there is text ambiguity in the user prompts, and searching for content-matching input for multi-modal. The third step is to construct a potential one-to-many instruction-answer pair based on the relationship between API functions.
Figure 7: Five types of ambiguity cases needed to be checked during the instruction-answer pair generation step.
8.1 Visualization of Dataset Construction Process
We show the whole process of collecting our dataset in Figure 6. The difficulty in collecting the benchmark is reflected in three aspects. (1) Task Classification: We restructure a hierarchical task classification system. This advancement rectifies the overly coarse-grained categorizations observed in the HuggingFace platform, ensuring that each API is endowed with a specific and unambiguous task label. (2) Annotation of Multimodal Input Instructions: Should an instruction reveal cues from modalities other than text, a precise coherence between the divulged information in the instruction and the content of the non-textual modality is paramount. In cases where no such cues are disclosed, we deem it essential to investigate the presence of five potential cases that are caused by textual ambiguity. And we show the visualization of five ambiguity types in Figure 7. Such scrutiny lays the groundwork for subsequent experiments aimed at validating whether information from other modalities aids large language models in making appropriate tool selections; (3) Instruction Matching: To identify all APIs meeting an instruction’s functional requirements, we must categorize APIs based on identical and similar functionalities respectively during collection. When pairing instructions for groups of APIs with analogous functions, it becomes crucial to discern whether a provided instruction strictly satisfies each API’s functional description. This determination will influence the number of APIs that can aptly respond as the correct answer to a given instruction.
8.2 Filtering Rules
After crawling APIs from HuggingFace, due to varying quality among these models, we implement stringent selection criteria for APIs. The specific filtering rules for this selection process are outlined as follows:
Low-Quality Model Card A significant number of APIs either lack model cards or contain overly simplistic ones, providing little helpful information. Additionally, some APIs have ceased maintenance or have been integrated into other APIs. Even multiple APIs share a single model card. These APIs pose challenges in organizing API documentation and require manual consultation of additional sources like papers and GitHub. While enriching the number of APIs, this approach introduces inaccuracies. Finally, we opted to remove such APIs from our dataset uniformly.
NSFW Content Risk For ethical and safety considerations, we remove APIs, particularly some of the text-to-image type, which pose risks of generating NSFW (Not Safe For Work) content. Such content is deemed inappropriate for all age groups and could compromise the dataset’s applicability in diverse settings.
License Restrictions We exclude models with restrictive licensing terms. An example is the Llama2  [34] model, which requires a formal application and approval process to download and use its model weights.
Identical APIs Numerous models on HuggingFace share identical structures and datasets, merely differing in implementation by various contributors. To avoid redundancy, we selectively retain only the most downloaded version of each such model, ensuring the dataset is more streamlined. For APIs with different model architectures but trained on the same dataset and having the same functions, as we mentioned in the main text, we select the top five APIs based on download statistics.
8.3 API Function Boundaries
Multi-task APIs Our dataset includes multi-task APIs like various large language models (Bloom  [36], Baichuan2  [40], Falcon  [22]), capable of numerous NLP tasks, some of which even fall outside HuggingFace’s classification system. For simplicity, we focus specifically on their text generation capabilities, aligning with HuggingFace’s functional categorization of such APIs. Additionally, we modify the descriptions of these APIs in our dataset to ensure a match between the functionalities reflected by the provided queries and those described in the API descriptions.
API Function Differentiation After identifying the specific subdomain of each API, we detail our approach for discerning whether APIs within the same subdomain possess identical or similar functions. For plain text input APIs, we can assess APIs based on the language and context of the datasets they utilize. For example, in sentiment analysis tasks, we discern various contexts by relying on the distinct domains of data sources, such as financial, legal, social media, and reviews, enabling us to classify nuanced functional differences among similar-function APIs. Similarly, for multimodal inputs, especially images, we refine API function categorization by analyzing dataset contexts. For instance, in Visual Question Answering tasks, datasets feature a range of images such as Diagrams, Charts, Documents, Infographics, and other general images, enabling nuanced API segmentation. The above examples provide merely a subset of our API classification rules. Due to the complex and case-by-case nature of the analysis, the complete set of rules is not listed here.
One-to-many Situation Identification Originating from each API’s functional description, our approach, driven by GPT-4, generates corresponding instructions that adhere strictly to the API’s functional boundaries. Hence, consideration is primarily given to other APIs with the same function in one-to-many scenarios.
8.4 Prompt Construction
Below, we list the detailed prompts used for generating instructions. The prompts fed to GPT-4 are composed of (1)the API call’s relevant information, including its name, a concise functional description, and training datasets if mentioned in the model card, (2)some sample instructions, and (3)additional requirements. The sample instructions are leveraged to foster format standardization and enhance the quality of GPT-4’s outputs. So we invite some expert annotators to craft some high-quality instructions in advance and then randomly choose two each time as exemplary inputs to GPT-4. We design two sets of prompt templates in our dataset, mindful of prompt ambiguity cases. For APIs with multimodal inputs susceptible to prompt ambiguity, we strive to balance the final selection of ten instructions, maintaining a 1:1 ratio between unambiguous and ambiguous instructions; the corresponding prompt template is shown in Figure 8 and Figure 9 respectively. We emphasize that ambiguity pairs in queries are those pairs with identical textual inputs but varied multimodal inputs.
[…]
Continued Description: This model is a base-sized ViT-based BLIP trained on the visual question answering task VQA2.0 dataset. The model’s input is an image and a text question, and the model’s output is a text answer. Prohibit Words: ”API, tools, model, VQA2.0” Output: Query1: I recently visited a new city and took some photographs. Based on this particular image, can you tell me what is the weather conditions like? Query2: I was at a photo exhibition and took this picture. Could you provide information about was this taken in a Latin American country judging from the dress? Query3: I was at the park and saw a dog playing. By analyzing this photo, can you tell if the frisbee it is using is hard or soft? Query4: My cousin went to a new food joint and ordered their specialty hot dog. By looking at this picture, can you determine if the hot dog is larger than a normal one? Query5: I attended a family gathering and took a group photo. By analyzing the photograph, how many kids can you see in the picture? Query6: Looking at this captivating portrait, I noticed an object in the individual’s hand. It seems to be a sweet treat. What is this person holding? Query7: I was at the beach and took a photo of a man skiing. By examining the image, can you tell if the wave is chasing him? Query8: I visited the local train station and captured an image of a passing train. Can you identify what color the train is from the picture? Query9: This cityscape presents an old-fashioned vehicle that stands out amidst the modern structures. I’d like to know more about this vehicle. Could you tell me its model and color?” Query20: Given the serene backdrop of sailboats and a calm sea, can you tell me how many boats are there? Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]
@@ 8.4 Prompt Construction
### User: You are an NLP task expert. Given an API, you need to generate 20 different user queries that utilize the API function, adhering to the following input and output format for each query: Input: API Name: This is the name of the API Function. Description: This is a detailed description of the model. Prohibit Words: These are words that are prohibited from appearing in the output. Output: Query1: This is an instruction that can utilize the API function. Query2: This is an instruction that can utilize the API function. Query3: This is an instruction that can utilize the API function. Query4: This is an instruction that can utilize the API function. Query5: This is an instruction that can utilize the API function. Query20: This is an instruction that can utilize the API function. Below are some examples: Example 1: Input: API Name: timm/resnet101.a1h_in1k Description: A 101 layers ResNet-B image classification model trained on ImageNet-1k. Prohibit Words: ”API, tools, model, ImageNet” Output: Query1: I’d value your input on this image’s classification. Query2: There’s an image in my possession, and I’m seeking a label for it. Query3: I need help in determining the label for this image. Can you help? Query4: Can you provide a label perspective for this image? Query5: How would you manage the categorization of this particular visual? Query6: There’s an intriguing image that I’ve come across, and I’m curious about its category. Query7: How would you classify this picture? Query8: I’d be grateful for your view on this image’s category. Query9: I’m searching for a label for this visual. Can you assist? Query20: What category seems plausible for this visual, in your view? Example 2: Input: API Name: lllyasviel/control_v11f1p_sd15_depth Description: This model is intended for control diffusion models by adding extra conditions. Trained with depth estimation, the condition image is an image with depth information, usually represented as a grayscale image, and the output is a new image. Prohibit Words: ”API, tools, model, depth, controlnet” Output: Query1: Use my image to depict a family picnic in a sunlit park with children playing. Query2: Can you reimagine my image as a soccer match in progress, with players and fans cheering? Query3: Create an adventure scene with pirates and treasure islands using my image. Query4: Can you make my image into a heartwarming scene of a mother duck leading her ducklings through a pond? To be continued
@@
Continued Query5: Craft a depiction of a magician performing tricks in front of an amazed audience from my image. Query6: Could you reshape this photograph into a bustling train station with intricate ironwork and arches? Query7: Transform this snapshot into a charming cottage nestled among rolling hills and wildflowers. Query8: I’m curious about how this picture would be reimagined as a bustling harbor with ships, cranes, and cargo containers. Query9: Use my image to depict a family picnic in a sunlit park with children playing. Query20: I’d love to see this image as a gardener carefully pruning roses in a blooming garden. Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response; 4. The queries should not convey or imply multimodal information. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]
@@
8.5 Visualization of Data Distributions

### v3

@@ 10.1 Results of Multiple Options Categorized by Input Modality
Continued
Example 2:
Input: API Name: Salesforce/blip-vqa-base Description: This model is a base-sized ViT-based BLIP trained on the visual question answering task VQA2.0 dataset. The model’s input is an image and a text question, and the model’s output is a text answer. Prohibit Words: ”API, tools, model, VQA2.0” Output: Query1: I recently visited a new city and took some photographs. Based on this particular image, can you tell me what is the weather conditions like? Query2: I was at a photo exhibition and took this picture. Could you provide information about was this taken in a Latin American country judging from the dress? Query3: I was at the park and saw a dog playing. By analyzing this photo, can you tell if the frisbee it is using is hard or soft? Query4: My cousin went to a new food joint and ordered their specialty hot dog. By looking at this picture, can you determine if the hot dog is larger than a normal one? Query5: I attended a family gathering and took a group photo. By analyzing the photograph, how many kids can you see in the picture? Query6: Looking at this captivating portrait, I noticed an object in the individual’s hand. It seems to be a sweet treat. What is this person holding? Query7: I was at the beach and took a photo of a man skiing. By examining the image, can you tell if the wave is chasing him? Query8: I visited the local train station and captured an image of a passing train. Can you identify what color the train is from the picture? Query9: This cityscape presents an old-fashioned vehicle that stands out amidst the modern structures. I’d like to know more about this vehicle. Could you tell me its model and color?” Query20: Given the serene backdrop of sailboats and a calm sea, can you tell me how many boats are there? Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]
@@
### User: You are an NLP task expert. Given an API, you need to generate 20 different user queries that utilize the API function, adhering to the following input and output format for each query: Input: API Name: This is the name of the API Function. Description: This is a detailed description of the model. Prohibit Words: These are words that are prohibited from appearing in the output. Output: Query1: This is an instruction that can utilize the API function. Query2: This is an instruction that can utilize the API function. Query3: This is an instruction that can utilize the API function. Query4: This is an instruction that can utilize the API function. Query5: This is an instruction that can utilize the API function. Query20: This is an instruction that can utilize the API function. Below are some examples: Example 1: Input: API Name: timm/resnet101.a1h_in1k Description: A 101 layers ResNet-B image classification model trained on ImageNet-1k. Prohibit Words: ”API, tools, model, ImageNet” Output: Query1: I’d value your input on this image’s classification. Query2: There’s an image in my possession, and I’m seeking a label for it. Query3: I need help in determining the label for this image. Can you help? Query4: Can you provide a label perspective for this image? Query5: How would you manage the categorization of this particular visual? Query6: There’s an intriguing image that I’ve come across, and I’m curious about its category. Query7: How would you classify this picture? Query8: I’d be grateful for your view on this image’s category. Query9: I’m searching for a label for this visual. Can you assist? Query20: What category seems plausible for this visual, in your view? Example 2: Input: API Name: lllyasviel/control_v11f1p_sd15_depth Description: This model is intended for control diffusion models by adding extra conditions. Trained with depth estimation, the condition image is an image with depth information, usually represented as a grayscale image, and the output is a new image. Prohibit Words: ”API, tools, model, depth, controlnet” To be continued
Continued
Output:
@@
Query1: Use my image to depict a family picnic in a sunlit park with children playing. Query2: Can you reimagine my image as a soccer match in progress, with players and fans cheering? Query3: Create an adventure scene with pirates and treasure islands using my image. Query4: Can you make my image into a heartwarming scene of a mother duck leading her ducklings through a pond? Query5: Craft a depiction of a magician performing tricks in front of an amazed audience from my image. Query6: Could you reshape this photograph into a bustling train station with intricate ironwork and arches? Query7: Transform this snapshot into a charming cottage nestled among rolling hills and wildflowers. Query8: I’m curious about how this picture would be reimagined as a bustling harbor with ships, cranes, and cargo containers. Query9: Use my image to depict a family picnic in a sunlit park with children playing. Query20: I’d love to see this image as a gardener carefully pruning roses in a blooming garden. Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response; 4. The queries should not convey or imply multimodal information. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]
@@
11 Dataset Visualization
11.1 Model Card Visualization
We illustrate 6 model card samples in Figure 10-15, which covers all the information mentioned in the main text. We list some samples from several different modalities and different functions.
11.2 Instruction-Answer Pairs Visualization
In Figure 16-18, we offer 7 cases sampled from the testing set. When we train our model, we input the conversation between humans and GPT in a conversation format, and the other attributes help us divide them into different sub-testing sets.
Figure 10: Model Card Visualization. (Sample 1)
Figure 11: Model Card Visualization. (Sample 2)
Figure 12: Model Card Visualization. (Sample 3)
Figure 13: Model Card Visualization. (Sample 4)
Figure 14: Model Card Visualization. (Sample 5)
Figure 15: Model Card Visualization. (Sample 6)
Figure 16: Instruction-Answer Pairs Visualization. (Sample 1-3)
Figure 17: Instruction-Answer Pairs Visualization. (Sample 4-5)
Figure 18: Instruction-Answer Pairs Visualization. (Sample 6-7)
11.3 Agent Output Visualization
We show four actual cases using our MLLM-Tool in Figure 19-22. Our system supports inputs from four modalities and leverages the predicted API to execute, finally getting customers’ expected results.
Figure 19: Agent Output Visualization. (Sample 1)
Figure 20: Agent Output Visualization. (Sample 2)
Figure 21: Agent Output Visualization. (Sample 3)
Figure 22: Agent Output Visualization. (Sample 4)